OpenStudy (anonymous):

what is a four or three step equation in which the answer is 4?

OpenStudy (anonymous):

3(6+5y) = 2(-5+4y) 18 + 15y = -10 + 8y 18 + 7y = -10 7y = -28 y = -4

OpenStudy (anonymous):

lets say our penultimate step will be 4x = 16 going back one step lets make it 4x + 1 = 16 +1 or 4x + 1 = 17 lets change 4x to 5x - x so we finally get 5x + 1 - x = 16

OpenStudy (anonymous):

sorry 5x + 1 - x = 17